OpenStudy (anonymous):

Solve the equation. X+1/6=6 A. -6 1/6 B. -5 5/6 C. 5 5/6 D. 6 1/6 @iGreen @uChezzy @perl @PaulaLovesSchool13 @chefducky please help

OpenStudy (anonymous):

@CausticSyndicalist

OpenStudy (anonymous):

x= 6-1/6 = 35/6 it's C.

OpenStudy (some.random.cool.kid):

im thinking c maybe.

OpenStudy (igreen):

@panchajanya @felavin @CausticSyndicalist Don't give answers, read the Code of Conduct. http://openstudy.com/code-of-conduct

OpenStudy (causticsyndicalist):

Ah okay. So shall I explain how to get it then?

OpenStudy (jhannybean):

subtract -1/6 from both sides x = 6 - 1/6 = 36-1/6 =?

OpenStudy (igreen):

\(x + \dfrac{1}{6} = 6\) Subtract \(\dfrac{1}{6}\) to both sides. \(x = 6 - \dfrac{1}{6}\)

OpenStudy (igreen):

@CausticSyndicalist No, we're supposed to guide people to the answer..you can't give an explanation with an answer..just give the explanation and let them get the answer on their own.

OpenStudy (causticsyndicalist):

Oh okay.

OpenStudy (anonymous):

okay okay please stop typing soo much I need to think

OpenStudy (igreen):

Well, if they want the answer in fraction form, you'll have to find the least common denominator to subtract, which would be 6 here. \(x = 6 - \dfrac{1}{6}\) \(x = \dfrac{6}{1} - \dfrac{1}{6}\) \(x = \dfrac{6 \times 6}{1 \times 6} - \dfrac{1 \times 1}{6 \times 1}\) \(x = \dfrac{36}{6} - \dfrac{1}{6}\) \(x = \dfrac{36-1}{6}\)

OpenStudy (anonymous):

i green i dont think that is the answer the asnwer choices are A. 2 B. 1 C.0

OpenStudy (anonymous):

THATS FOR MY QUESTION THAT U ANSWERD

OpenStudy (igreen):

LOL

OpenStudy (igreen):

@sammixboo You should use \dfrac instead of \frac \(\dfrac{1}{2}\) - \dfrac \(\frac{1}{2}\) - \frac

sammixboo (sammixboo):

Oooo I see

OpenStudy (igreen):

Plus you shouldn't give the full answer..